what does 5/3w=-10 equal?

Found 2 solutions by Alan3354, richard1234:
Answer by Alan3354(69443)

what does 5/3w=-10 equal?
--------
It equals -10, and it equals 5/3w.
That's what you said.

Answer by richard1234(7193)

The question "what does 5/3w = -10" equal" doesn't make sense. Instead, you would rather say "5/3w = -10, what does w equal?"

Additionally, are you saying

, or

?

If then you can multiply both sides by 3/5, yielding w = -6. Or, if then cross-multiplying, -30w = 5, w = -1/6.